Please Please Help Me!!!
I am stuck on one question that I have probably trying to figure out for about an hour I understand what I'm supposes to do but just can get the right answer???
So I would really appreciate if someone would kindly help me.
This is the question
Solve by substitution
x-5y=34...(1)
3x+4y=-12..(2)
This is how I'm doing it
-5y=(34-x).....(1)
3x+4(34-x)=-12..(2)
3x-4(-34+x)=-12
3x+136-4x=-12
-1x +136=-12
-1x=-12+136
-1x=124
-1 -1
x=-124
& I know this is not the right answer for (x)I've even tried to change my signs & even substitute the other equation
Please Help Me!!!
It would really mean a lot.
Thank you

Answer by edjones(8007)

-5y=(34-x).....(1) OK so far.
y=-(34-x)/5
=(x-34)/5 is what you should have done. you substituted 34-x for y, but -5y=34-x, not y.
You could have solved the problem by substituting for y but you have an unwieldy
fraction to substitute.
Much better is to substitute for x:
x-5y=34 (1)
x=5y+34
3x+4y=-12
3(5y+34)+4y=-12
15y+102+4y=-12
19y=-114
y=-6
x=4
.
Keep up the struggle. You WILL get it.
.
Ed